# DiaPer: End-to-End Neural Diarization with Perceiver-Based Attractors

### v2

@@ Abstract
Until recently, the field of speaker diarization was dominated by cascaded systems. Due to their limitations, mainly regarding overlapped speech and cumbersome pipelines, end-to-end models have gained great popularity lately. One of the most successful models is end-to-end neural diarization with encoder-decoder based attractors (EEND-EDA). In this work, we replace the EDA module with a Perceiver-based one and show its advantages over EEND-EDA; namely obtaining better performance on the largely studied Callhome dataset, finding the quantity of speakers in a conversation more accurately, and running inference on almost half of the time on long recordings. Furthermore, when exhaustively compared with other methods, our model, DiaPer, reaches remarkable performance with a very lightweight design. Besides, we perform comparisons with other works and a cascaded baseline across more than ten public wide-band datasets. Together with this publication, we release the code of DiaPer as well as models trained on public and free data.
@@ I Introduction
Since the appearance of end-to-end models, the ecosystem has changed substantially with new approaches constantly appearing [4]. Neural-based diarization models can be separated into different categories: single-stage systems, which comprise only one model, and two-stage systems, which have two steps where one is a variant of end-to-end model and the other is either based on clustering or on another model. Single-stage systems, such as end-to-end neural diarization (EEND) [5], where diarization is modeled as per-speaker per-frame binary classification, are trained directly for the task. While the training can be done in different steps, the inference is performed in a single stage. These methods face difficulties in recordings with several speakers [6]. Two-stage systems can be separated into different classes. Models such as target speaker voice activity detection [7] are trained in an end-to-end manner but make use of an initialization provided by an existing (usually cascaded) model which has to be run priorly at inference time. Other two-stage systems run EEND on short segments (where few speakers are expected) and then perform clustering to join the decisions on short segments. They are known as EEND vector clustering (EEND-VC) and different variants have been proposed [8, 9, 10]. These approaches present advantages in dealing with several speakers (potentially an unlimited number of them) while having an edge over clustering-based methods on dealing with overlapped speech segments as EEND models usually do. This categorization is, however, not strict. Some systems do not exactly qualify as “single” or “two” stage as they have a single stage but include some iterative procedure [11, 12].
@@ III The Model
Figure 1: DiaPer diagram.
@@
In other words, the frame encoder is in charge of transforming the initial input features into deeper and more contextualized representations from which (a) the attractors will be estimated, and (b) the frame-wise activation of each speaker will be determined. Several encoder layers are used to extract such representations and, in a similar way as presented in [27], each layer also includes frame-speaker activities conditioning. As shown in Figure 2, intermediate attractors are calculated given the frame embeddings of each frame encoder layer. The intermediate attractors are then weighted by intermediate frame activities and transformed into the frame embedding space to produce the conditioning.
@@
where $PercDec$ is the Perceiver-based attractor decoder, $\textbf{W}_{c}\in\mathbb{R}^{D\times D}$ is a learnable parameter that weights the effect of the intermediate attractors on the frame embeddings.
@@
The decoder makes use of a chain of a few Perceiver blocks as depicted in Figure 3. The set of learnable latents is transformed by each block utilizing the frame embeddings as keys and values. One could have an equal amount of latents and attractors, in which case the latents are an initial representation transformed by the blocks to obtain the attractors. In practice, we observed that this leads to instability in the training and that obtaining the attractors as the linear combination of a larger set of (transformed) latents performed better. More formally,
@@ IV-A2 Evaluation data
Besides telephone conversations, we compared the models on a variety of wide-band datasets. As the models we evaluate are trained on single-channel data, when the datasets contain microphone array data, we mix all channels in the microphone array (far-field) or headsets (near-field) [^5] It should be noted that other works in Table XII might have processed the channels differently so the acoustic inputs for the recordings might differ.. Training sets (or development, if train sets are not available) are utilized for fine-tuning. The databases considered are:
@@ IV-B Models
Following standard practice with EEND models, the training scheme consists in training the model first on synthetic training data and then performing fine-tuning (FT) using a small development set of real data of the same domain as the test set. In the experiments with more than two speakers, a model initially trained on synthetic data with two speakers per recording is adapted to a synthetic set with a variable number of speakers and finally fine-tuned to a development set.
@@ V-A Selection of parameters
In order to shed some light on the influence of different aspects of the architecture in DiaPer, we present first a comparison of the performance when varying some key elements. We start from the best configuration we found, namely: 3 Perceiver blocks in the attractor decoder, 128 latents, 4 self-attention layers in the frame-encoder and 128-dimensional latents, frame embeddings and attractors. This configuration is marked with a gray background in the comparisons. The models are trained on 2-speaker SC and no FT is applied.
Table III shows the impact of the number of Perceiver blocks in the attractor decoder. Out of the configurations explored, having 3 blocks presents the best performance.
TABLE III: Comparison on CH1-2spk when varying the number of Perceiver blocks in the attractor decoder.
@@
Table IV shows how the number of latents can affect the performance. Differences are small for all amounts equal to or below 256, even with as few as 8. Nevertheless, given that the number of parameters is very similar for any configuration, we keep 128 latents as having more could ease the task when more speakers appear in a recording.
TABLE IV: Comparison on CH1-2spk when varying the number of latents.
@@
Table V presents a comparison when varying the number of layers in the frame encoder. Standard SA-EEND and EEND-EDA use 4 and some works have used 6 layers. In the case of DiaPer, we do not observe large differences in the performance and obtain the best performance with 4.
TABLE V: Comparison on CH1-2spk when varying the number of layers in frame encoder.
@@
Finally, Table VI shows the impact of the model dimensions on the performance. Increasing the dimensionality of latents, frame embeddings and attractors further than 128 does not show improvements in terms of DER but increases the number of model parameters significantly. Figure 4 shows performance throughout the epochs for the development set. It is clear how more dimensions allow for a faster convergence; however, more than 128 do not provide more gains in terms of final performance. In addition, more dimensions make the training less stable: using 512 would always lead to instability. Configurations with less than 128 dimensions (64 and 32) can improve further and after 200 epochs reduce the DER by about 1 point but still with worse final results than other configurations. These findings show that reasonable performances can be achieved even with more lightweight versions of DiaPer.
@@
TABLE VI: Comparison on CH1-2spk when varying the model dimension (latents, frame embeddings and attractors).
@@ V-B Ablation analysis
Different decisions were made when developing DiaPer and some have a big impact on the performance. Table VII presents a comparison of DiaPer in the best configuration shown above and when removing some of the operations performed during training. The first one refers to the normalization of the loss by the reference quantity of speakers, as shown in Eq. 19. DiaPer always outputs $A$ attractors and the loss is calculated for all of them, even if only training with 2-speakers SC. If the loss is not normalized by the amount of speakers, the model tends to find less speech, increasing the missed speech rate considerably.
@@
Finally, the attention normalization in the cross-attention calculations inside the Perceiver blocks is performed across latents in DiaPer. If done across time, as it is usually done, slightly worsens the performance. We have also explored using across-time normalization in half of the heads and across-latents in the other half but the performance was not better than using across-latents in all heads.
TABLE VII: DER (%) on CH1-2spk with different ablation comparisons.
@@ V-C Two-speaker telephone conversations
Figure 6 presents a comparison between EEND-EDA and DiaPer inference times. Although DiaPer is slower for very short recordings, it can run considerably faster when processing several-minute recordings. This speed-up is given only by the Perceiver-based attractor decoder (instead of the LSTM-based of EEND-EDA) since the rest of the model is the same. Even more, these results correspond to an input downsampling factor of 10 but if more precision was used, the length of the frame-embedding sequences increases which would show further advantage for DiaPer for the same recording lengths. Notably, EEND-EDA has 6.4 million parameters while DiaPer has only 4.6 million showing that the model not only runs faster when processing long sequences but also makes more efficient use of the parameters.
@@
Table VIII presents an exhaustive comparison with all competitive systems at the time of publication under the same conditions: all speech is evaluated and no oracle information is used. Data refers to the number of hours of data for supervision. For end-to-end models, it can be real or synthetic data and for the clustering-based baseline, it consists of all data used to train the x-vector extractor, VAD and OSD. Methods are divided into groups depending on if they are single or two-stage. Even though DiaPer does not present the best performance among all approaches, it reaches competitive results with fewer parameters and even without FT.
TABLE VIII: DER (%) comparison on CH2-2spk with other methods. For our results, we selected the model with the best performance on CH1 out of the 5 runs. Type can be clustering (C), 1-stage (1-S), or 2-stage (2-S) system. (I) stands for iterative, meaning there is an iterative process at inference time.
@@
DiaPer 1-S 4.6 2.5 $8.05$ 7.517.517.517.51121212It is worth mentioning that out of the 5 runs, the best DER on Part 2 was 7.38 but that did not correspond to the lowest DER on Part 1. Analogously, for EEND-EDA it was 7.78.
@@ V-D Multiple-speakers telephone conversations
Figure 7 presents the comparison for recordings with multiple amounts of speakers where EEND-EDA and DiaPer are trained on the same data. Once again, DiaPer presents significant advantages over EEND-EDA both before and after fine-tuning to the development set. Table IX shows the DER for different numbers of speakers per conversation where gains are observed in almost all cases. The largest differences are for recordings with more speakers, suggesting the superiority of DiaPer in handling such situations.
@@
TABLE IX: DER (%) comparison on CH2. For each method (EEND-EDA and DiaPer), selecting the best model on CH1 out of the 5 runs.
@@
Finally, Table X shows the comparison of DER components. It can be observed that without fine-tuning DiaPer does not improve the confusion error of EEND-EDA but rather missed and false alarm (FA) speech. A closer look at the inherent VAD and OSD performances of the two models allows us to see that DiaPer improves considerably the OSD recall with similar OSD precision. Therefore, most of the improvement is related to more accurate overlapped speech detection. Nevertheless, it should be pointed out that precision and recall slightly above 50% are still very low. There is clearly large room for improving the performance in this aspect.
TABLE X: Comparison on CH2. For each method, selecting the best model on CH1 out of the 5 runs. DER and its three components and precision and recall for VAD and OSD performance.
@@
Finally, Table XI presents comparisons with other publications on Callhome Part 2 using all recordings. Again, all speech is evaluated and no oracle information is used. For these comparisons, we utilize one of the models trained seeing SC up to 7 speakers (since Callhome does not contain recordings with more speakers). Results show that even if DiaPer has a competitive performance, many methods can reach considerably better results. The main advantage of DiaPer is its lightweight nature, having the least number of parameters in comparison with all other methods. Exploring larger versions of DiaPer (i.e. increasing the model dimension) which could lead to better performance in multi-speaker scenarios is left for future research.
@@
TABLE XI: DER comparison on CH2 with other methods. For our results, we selected the model with the best performance on CH1 out of the 5 runs. Type can be clustering (C), 1-stage (1-S) or 2-stage (2-S) system. (I) stands for iterative, meaning there is an iterative process at inference time.
@@
DiaPer 1-S 4.6 15 14.86 13.6013.6013.6013.60131313It is worth mentioning that out of the 5 runs, the best DER on Part 2 was 13.16 but that did not correspond to the lowest DER on Part 1.
@@
DiaPer 1-S 4.6 2.5 27.84 24.1624.1624.1624.16141414It is worth mentioning that out of the 5 runs, the best DER on Part 2 was 23.81 but that did not correspond to the lowest DER on Part 1.
TABLE XII: DER (%) comparison on a variety of test sets. Overlaps are evaluated and oracle VAD is NOT used. SR stands for sampling rate. Underlined results denote single systems and $\overline{\mbox{overlined}}$ results correspond to fusions or more complex models.
System
@@
VAD+VBx+OSD 8 14.46 29.44 22.68 34.14 22.23 84.01 27.86 20.49 56.19 38.09 18.81 18.33 6.69
DiaPer 8 49.29 45.40 33.70 54.69 41.31 78.51 49.92 38.24 64.33 19.11 34.62 32.61 32.37
DiaPer+FT 8 42.66 31.60 28.93 50.49 36.41 68.54 34.10 24.08 45.09 14.69 17.99 20.90 31.63
VAD+VBx+OSD 16 15.84 28.84 22.59 34.61 22.42 70.42 26.67 20.28 49.22 35.60 16.86 18.19 6.12
DiaPer 16 48.21 38.67 28.19 57.07 36.36 78.25 43.75 34.21 48.26 21.03 35.69 38.05 23.20
DiaPer+FT 16 41.43 32.60 27.82 49.75 32.94 70.77 32.97 24.12 Overfit 13.41 15.46 21.11 Overfit
DiaPer (20att) 16 47.86 34.35 23.90 52.29 35.08 77.51 44.51 34.82 43.37 18.51 25.07 32.08 22.10
DiaPer (20att)+FT 16 31.30 26.27 24.44 50.97 30.49 69.94 31.23 22.77 Overfit 10.99 14.59 18.69 Overfit
Best 16.76[80] 23.8[10] 22.2[10] 18.0[74] $\overline{\mbox{32.46}}$[81] 26.4[9] 17.32[82] $\overline{\mbox{22.36}}$[83] $\overline{\mbox{7.27}}$[83] 21.96[63] 22.2[74] $\overline{\mbox{4.0}}$[84]
published 14.0[10] 23.3[74] 22.0[74] 16.95[82] $\overline{\mbox{27.25}}$[83] 26.88[85] $\overline{\mbox{16.94}}$[86] $\overline{\mbox{22.04}}$[87] $\overline{\mbox{6.14}}$[87] 33.6[88] 19.90[64] $\overline{\mbox{4.39}}$[89]
results 13.2[74] 23.5[90] 19.53[82] 13.00[29] $\overline{\mbox{25.11}}$[87] 24.64[29] $\overline{\mbox{16.76}}$[82] $\overline{\mbox{16.36}}$[87] $\overline{\mbox{5.65}}$[81] 16.0[74] 14.37[25] $\overline{\mbox{4.35}}$[91]

### v3

@@ Abstract
Until recently, the field of speaker diarization was dominated by cascaded systems. Due to their limitations, mainly regarding overlapped speech and cumbersome pipelines, end-to-end models have gained great popularity lately. One of the most successful models is end-to-end neural diarization with encoder-decoder based attractors (EEND-EDA). In this work, we replace the EDA module with a Perceiver-based one and show its advantages over EEND-EDA; namely obtaining better performance on the largely studied Callhome dataset, finding the quantity of speakers in a conversation more accurately, and faster inference time. Furthermore, when exhaustively compared with other methods, our model, DiaPer, reaches remarkable performance with a very lightweight design. Besides, we perform comparisons with other works and a cascaded baseline across more than ten public wide-band datasets. Together with this publication, we release the code of DiaPer as well as models trained on public and free data.
@@ I Introduction
Since the appearance of end-to-end models, the ecosystem has changed substantially with new approaches constantly appearing [4]. Neural-based diarization models can be separated into different categories: single-stage systems, which comprise only one model, and two-stage systems, which have two steps where one is a variant of end-to-end model and the other is either based on clustering or on another model. Single-stage systems, such as end-to-end neural diarization (EEND) [5], where diarization is modeled as per-speaker per-frame binary classification, are trained directly for the task. While the training can be done in different steps (training with 2-speaker simulated data, then adapting to data with variable number of speakers and finally fine-tuning to in-domain data), the inference is performed in a single stage. These methods face difficulties in recordings with several speakers [6]. Two-stage systems can be separated into different classes. Models such as target speaker voice activity detection [7] are trained in an end-to-end manner but make use of an initialization provided by an existing (usually cascaded) model which has to be run priorly at inference time. Other two-stage systems run EEND on short segments (where few speakers are expected) and then perform clustering to join the decisions on short segments. They are known as EEND vector clustering (EEND-VC) and different variants have been proposed [8, 9, 10]. These approaches present advantages in dealing with several speakers (potentially an unlimited number of them) while having an edge over clustering-based methods on dealing with overlapped speech segments as EEND models usually do. This categorization is, however, not strict. Some systems do not exactly qualify as “single” or “two” stage as they have a single stage but include some iterative procedure [11, 12].
@@ III The Model
Figure 1: DiaPer diagram. $\sigma$ refers to the sigmoid function and the circles with crosses mean dot-product between the vectors.
@@
In other words, the frame encoder is in charge of transforming the initial input features into deeper and more contextualized representations from which (a) the attractors will be estimated, and (b) the frame-wise activation of each speaker will be determined. Several encoder layers are used to extract such representations and, in a similar way as presented in [27], each layer also includes frame-speaker activities conditioning. As shown in Figure 2, intermediate attractors are calculated given the frame embeddings of each frame encoder layer. The intermediate attractors are then weighted by intermediate frame activities and transformed into the frame embedding space to produce the conditioning. While in EEND-EDA the input frame embeddings are directly processed to obtain attractors, attractors obtained with an attention mechanism need queries to be compatible with keys. The intermediate loss ensures that they match at different encoder layers, thus easing the compatibility at the end of the frame encoder. The attractors are always calculated with the same Perceiver-based decoder, i.e. the parameters are shared for all the intermediate attractors.
@@
where $PercDec$ is the Perceiver-based attractor decoder, $\textbf{W}_{c}\in\mathbb{R}^{D\times D}$ is a learnable parameter that weights the effect of the intermediate attractors on the frame embeddings. The application of the conditioning mechanism allows the intermediate frame embeddings to be contextualized given the attractors [^2] This mechanism could also be understood as a cross-attention operation where the frame embeddings function as queries and the attractors as keys and values, and the attention weights are given by the frame-speaker activities.. We utilize the term “conditioning” to be consistent with [27].
@@
The decoder makes use of a chain of a few Perceiver blocks as depicted in Figure 3. The set of learnable latents is transformed by each block utilizing the frame embeddings as keys and values. The latents are randomly initialized before starting training and learned during that process. Then, they are fixed at inference time and transformed with the successive application of the Perceiver blocks to be adapted for the given recording.
One could have an equal number of latents and attractors, in which case the latents are an initial representation transformed by the blocks to obtain the attractors. In practice, we observed that this leads to instability in the training and that obtaining the attractors as the linear combination of a larger set of (transformed) latents performed better. More formally,
@@ IV-A2 Evaluation data
Besides telephone conversations, we compared the models on a variety of wide-band datasets. As the models we evaluate are trained on single-channel data, when the datasets contain microphone array data, we mix all channels in the microphone array (far-field) or headsets (near-field) [^6] It should be noted that other works in Table XIII might have processed the channels differently so the acoustic inputs for the recordings might differ.. Training sets (or development, if train sets are not available) are utilized for fine-tuning. The databases considered are:
@@ IV-B Models
Following standard practice with EEND models, the training scheme consists in training the model first on synthetic training data and then performing fine-tuning (FT) using a small development set of real data of the same domain as the test set. In the experiments with more than two speakers, a model initially trained on synthetic data with two speakers per recording is adapted to a synthetic set with a variable number of speakers and finally fine-tuned to a development set. We explored training directly on a set with a variable number of speakers and, as in our previous work [34], we observed that for the same training time the model would not reach the same performance as training on 2-speaker SC and then adapting to a variable number of speakers. While this does not mean that both approaches cannot reach the same performance, training on a large set with variable number of speakers is costly. From a practical perspective, training first on a 2-speaker set and then adapting to more speakers results in shorter training times. Nevertheless, other curriculum learning strategies could result in more overall efficient training pipelines for Diaper and all other EEND systems, which is something yet to be explored by the community.
@@ IV-C Training
TABLE III: Information about training steps. “*” in the fourth column refers to different numbers of epochs for different experiments. The sixth column refers to the maximum number of epochs but different experiments used different (not always the last) epochs. The warm-up in Noam was always carried out for 200k steps.
SR Step ID Init. (# ep.) Set # max. ep. LR Related results
8 kHz Training (2-speaker SC) A None SC2 8 kHz 100 Noam Tables IV, V, VI, VII, VIII, IX
and Figures 4, 5
Adaptation (multi-speaker SC) B A* SC2-7 8 kHz 75 Noam Figures 8, 9 and Tables X, XI
C A* SC1-10 8 kHz 100 Noam Figures 9, 10, Table XIII (2)
Fine-tuning (with in-domain data) D A* CH1 2 speakers 20 Adam $10^{-5}$ Figure 5 and Table IX
E A* DIHARD 3 CTS dev 20 Adam $10^{-4}$ Figure 5
F B CH1 20 Adam $10^{-4}$ Figure 8 and Tables X, XI, XII
16 kHz Training (2-speaker SC) G None SC2 16 kHz 100 Noam None
Adaptation (multi-speaker SC) H G (90-100) SC1-10 16 kHz 100 Noam Table XIII (5) and (7)
Fine-tuning (with in-domain data) I H (90-100) Various Various Adam $10^{-6}$ Table XIII (6) and (8)
Fine-tuning (with compound set) J H (90-100) Compound 400 Adam $10^{-5}$ Table XIII (9), (10) and (11)
Fine-tuning (with in-domain data) K J (390-400) Various Various Adam $10^{-6}$ Table XIII (12)
@@ V-A Selection of parameters
In order to shed some light on the influence of different aspects of the architecture in DiaPer, we present first a comparison of the performance when varying some key elements. We start from the best configuration we found, namely: 3 Perceiver blocks in the attractor decoder, 128 latents, 4 self-attention layers in the frame-encoder and 128-dimensional latents, frame embeddings and attractors. This configuration is marked with a gray background in the comparisons. The models are trained on 2-speaker SC and no FT is applied. We also considered different numbers of attractors: 5, 10, and 20 but the performance was the same for the 2-speaker scenario. All the experiments in Sections V-A,V-B,V-C,V-D had models with 10 attractors which is an upper bound on the expected number of speakers in a recording.
Table IV shows the impact of the number of Perceiver blocks in the attractor decoder. Out of the configurations explored, having 3 blocks presents the best performance.
TABLE IV: Comparison on CH1-2spk when varying the number of Perceiver blocks in the attractor decoder.
@@
Table V shows how the number of latents can affect the performance. Differences are small for all amounts equal to or below 256, even with as few as 8. Nevertheless, given that the number of parameters is very similar for any configuration, we keep 128 latents as having more could ease the task when more speakers appear in a recording.
TABLE V: Comparison on CH1-2spk when varying the number of latents.
@@
Table VI presents a comparison when varying the number of layers in the frame encoder. Standard SA-EEND and EEND-EDA use 4 and some works have used 6 layers. In the case of DiaPer, we do not observe large differences in the performance and obtain the best performance with 4.
TABLE VI: Comparison on CH1-2spk when varying the number of layers in frame encoder.
@@
Finally, Table VII shows the impact of the model dimensions on the performance. Increasing the dimensionality of latents, frame embeddings and attractors further than 128 does not show improvements in terms of DER but increases the number of model parameters significantly. Figure 4 shows performance throughout the epochs for the development set. It is clear how more dimensions allow for a faster convergence; however, more than 128 do not provide more gains in terms of final performance. In addition, more dimensions make the training less stable: using 512 would always lead to instability. Configurations with less than 128 dimensions (64 and 32) can improve further and after 200 epochs reduce the DER by about 1 point but still with worse final results than other configurations. These findings show that reasonable performances can be achieved even with more lightweight versions of DiaPer.
@@
TABLE VII: Comparison on CH1-2spk when varying the model dimension (latents, frame embeddings and attractors).
@@ V-B Ablation analysis
Different decisions were made when developing DiaPer and some have a big impact on the performance. Table VIII presents a comparison of DiaPer in the best configuration shown above and when removing some of the operations performed during training. The first one refers to the normalization of the loss by the reference quantity of speakers, as shown in Eq. 19. DiaPer always outputs $A$ attractors and the loss is calculated for all of them, even if only training with 2-speakers SC. If the loss is not normalized by the amount of speakers, the model tends to find less speech, increasing the missed speech rate considerably.
@@
The attention normalization in the cross-attention calculations inside the Perceiver blocks is performed across latents in DiaPer. If done across time, as it is usually done, slightly worsens the performance. We have also explored using across-time normalization in half of the heads and across-latents in the other half but the performance was not better than using across-latents in all heads.
Finally, we also explored removing the entropy loss (Eq. 18). While the performance is only slightly lower without the loss, the effect might be larger when handling many speakers. The performance on the whole CH1 set was worse in this case, even if the models were trained only with 2-speaker SC.
TABLE VIII: DER (%) on CH1-2spk with different ablation comparisons.
@@
  • Use a linear layer to transform latents into attractors instead of a simple linear combination (learnable) matrix (worse performance).
@@ V-C Two-speaker telephone conversations
Figure 7: Average inference memory footprint for both EEND-EDA and DiaPer models for recordings from 1 minute to 1 hour running 5 times each inference with a downsampling factor of 10. Experiments ran on Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz.
Figure 6 presents a comparison between the standard EEND-EDA baseline and DiaPer inference times. Although DiaPer is slower for very short recordings, it can run faster than the standard EEND-EDA when processing several-minute recordings. This speed-up is mainly given by the more light-weight nature of DiaPer which results in a faster frame encoder processing, which dominates the computation time versus the attractor decoding in both models. To have a fair comparison, “EEND-EDA small” denotes a version of EEND-EDA where the model dimension matches that of DiaPer in its best configuration (128-dimensional frame embeddings and attractors). This corresponds to using the same frame encoder in both models and we can see that DiaPer is slightly slower due to more computations in the attractor decoder. It should be noted that EEND-EDA small performs slightly worse than EEND-EDA in terms of DER. This was not the case for DiaPer and with the smaller configuration, we are able to obtain better DER performance than EEND-EDA and run faster.
Figure 7 presents the memory footprint of the models. Note that both EEND-EDA and DiaPer share the same self-attention-based frame encoder architecture which is the main source of memory consumption. Therefore, the memory footprint is a direct function of the sequence length, not the number of parameters of the models. It should be pointed out that these models have very high requirements for long recordings. There is certainly room for improvement regarding this aspect to make end-to-end models more memory efficient.
Table IX presents an exhaustive comparison with all competitive systems at the time of publication under the same conditions: all speech is evaluated and no oracle information is used. Data refers to the number of hours of data for supervision. For end-to-end models, it can be real or synthetic data and for the clustering-based baseline, it consists of all data used to train the x-vector extractor, VAD and OSD. Methods are divided into groups depending on if they are single or two-stage. Even though DiaPer does not present the best performance among all approaches, it reaches competitive results with fewer parameters and even without FT.
TABLE IX: DER (%) comparison on CH2-2spk with other methods. For our results, we selected the model with the best performance on CH1 out of the 5 runs. Type can be clustering (C), 1-stage (1-S), or 2-stage (2-S) system. (I) stands for iterative, meaning there is an iterative process at inference time.
@@
DiaPer 1-S 4.6 2.5 $8.05$ 7.517.517.517.51131313It is worth mentioning that out of the 5 runs, the best DER on Part 2 was 7.38 but that did not correspond to the lowest DER on Part 1. Analogously, for EEND-EDA it was 7.78.
@@ V-D Multiple-speakers telephone conversations
Figure 8 presents the comparison for recordings with multiple amounts of speakers where EEND-EDA and DiaPer are trained on the same data. Once again, DiaPer presents significant advantages over EEND-EDA both before and after fine-tuning to the development set. Table X shows the DER for different numbers of speakers per conversation where gains are observed in almost all cases. The largest differences are for recordings with more speakers, suggesting the superiority of DiaPer in handling such situations. However, it should be noted that there are only 3 files with 6 speakers and improvements in only one file can affect the results considerably, as is the case here.
@@
TABLE X: DER (%) comparison on CH2. For each method (EEND-EDA and DiaPer), selecting the best model on CH1 out of the 5 runs.
@@
Table XI shows the comparison of DER components. It can be observed that without fine-tuning DiaPer does not improve the confusion error of EEND-EDA but rather missed and false alarm (FA) speech. A closer look at the inherent VAD and OSD performances of the two models allows us to see that DiaPer improves considerably the OSD recall with similar OSD precision. Therefore, most of the improvement is related to more accurate overlapped speech detection. Nevertheless, it should be pointed out that precision and recall slightly above 50% are still very low. There is clearly large room for improving the performance in this aspect.
TABLE XI: Comparison on CH2. For each method, selecting the best model on CH1 out of the 5 runs. DER and its three components and precision and recall for VAD and OSD performance.
@@
Finally, Table XII presents comparisons with other publications on Callhome Part 2 using all recordings. Again, all speech is evaluated and no oracle information is used. For these comparisons, we utilize one of the models trained seeing SC up to 7 speakers (since Callhome does not contain recordings with more speakers). Results show that even if DiaPer has a competitive performance, many methods can reach considerably better results. The main advantage of DiaPer is its lightweight nature, having the least number of parameters in comparison with all other methods. Exploring larger versions of DiaPer (i.e. increasing the model dimension) which could lead to better performance in multi-speaker scenarios is left for future research.
@@
TABLE XII: DER comparison on CH2 with other methods. For our results, we selected the model with the best performance on CH1 out of the 5 runs. Type can be clustering (C), 1-stage (1-S) or 2-stage (2-S) system. (I) stands for iterative, meaning there is an iterative process at inference time.
@@
DiaPer 1-S 4.6 15 14.86 13.6013.6013.6013.60141414It is worth mentioning that out of the 5 runs, the best DER on Part 2 was 13.16 but that did not correspond to the lowest DER on Part 1.
@@
DiaPer 1-S 4.6 2.5 27.84 24.1624.1624.1624.16151515It is worth mentioning that out of the 5 runs, the best DER on Part 2 was 23.81 but that did not correspond to the lowest DER on Part 1.
TABLE XIII: DER (%) comparison on a variety of test sets. Overlaps are evaluated and oracle VAD is NOT used. SR stands for sampling rate and “Cp.” refers to the compound training set. Results with “*” are worse on the test set after fine-tuning but the decision was made on the development set, for which there were improvements. “Best published results” refers to the best three reported results at the time of writing. Underlined results denote single systems and $\overline{\mbox{overlined}}$ results correspond to fusions or more complex models.
ID System
@@
(1) VAD+VBx+OSD 8 14.5 29.4 22.7 34.1 22.2 84.0 27.9 20.5 56.2 38.1 18.8 18.3 6.7
(2) DiaPer (10att) 8 49.3 45.4 33.7 54.7 41.3 78.5 49.9 38.2 64.3 19.1 34.6 32.6 32.4
(3) DiaPer+FT 8 42.7 31.6 28.9 50.5 36.4 68.5 34.1 24.1 45.1 14.7 18.0 20.9 31.6
(4) VAD+VBx+OSD 16 15.8 28.8 22.6 34.6 22.4 70.4 26.7 20.3 49.2 35.6 16.9 18.2 6.1
(5) DiaPer (10att) 16 48.2 38.7 28.2 57.1 36.4 78.3 43.8 34.2 48.3 21.0 35.7 38.1 23.2
(6) (5) + FT 1m 16 41.4 32.6 27.8 49.8 32.9 70.8 33.0 24.1 Overfit 13.4 15.5 21.1 Overfit
(7) DiaPer (20att) 16 47.9 34.4 23.9 52.3 35.1 77.5 44.5 34.8 43.4 18.5 25.1 32.1 22.1
(8) (7) + FT 1m 16 31.3 26.3 24.4* 51.0 30.5 69.9 31.2 22.8 Overfit 11.0 14.6 18.7 Overfit
(9) (7) + FT Cp. 1m 16 39.5 27.1 23.4 53.1 33.0 67.7 34.1 25.7 44.7 14.9 14.9 21.7 26.8
(10) (7) + FT Cp. 5m 16 29.4 20.7 18.2 45.3 23.9 62.6 29.1 21.1 39.6 21.9 15.3 16.2 19.6
(11) (7) + FT Cp. 10m 16 29.0 21.3 17.8 40.7 24.6 61.1 29.9 21.8 36.1 24.0 16.0 16.1 19.1
(12) (7-9) + FT 16 28.8 20.2 17.6 37.5 29.1* 61.6* 27.7 20.3 33.8 12.5 13.4 15.7 18.2
Best 16.8[80] 23.8[10] 22.2[10] 18.0[74] $\overline{\mbox{32.5}}$[81] 26.4[9] 17.3[82] $\overline{\mbox{22.4}}$[83] $\overline{\mbox{7.3}}$[83] 22.0[63] 22.2[74] $\overline{\mbox{4.0}}$[84]
published 14.0[10] 23.3[74] 22.0[74] 17.0[82] $\overline{\mbox{27.3}}$[83] 26.9[85] $\overline{\mbox{16.9}}$[86] $\overline{\mbox{22.0}}$[87] $\overline{\mbox{6.1}}$[87] 33.6[88] 19.9[64] $\overline{\mbox{4.4}}$[89]
results 13.2[74] 23.5[90] 19.5[82] 13.0[29] $\overline{\mbox{25.1}}$[87] 24.6[29] $\overline{\mbox{16.8}}$[82] $\overline{\mbox{16.4}}$[87] $\overline{\mbox{5.7}}$[81] 16.0[74] 14.4[25] $\overline{\mbox{4.4}}$[91]